OpenStudy (anonymous):

When solving the system x + 2y = 9 and x - y = 6, which expression could be substituted for x in the first equation?

OpenStudy (anonymous):

solve for x in the 2nd equation x = 6+y

OpenStudy (anonymous):

How do you solve it though?

OpenStudy (anonymous):

x - y = 6 add y to both sides x = 6-y

OpenStudy (anonymous):

sorry x = 6+y

OpenStudy (anonymous):

x + 2y = 9 and x -y = 6, so x = y + 6 substitute that for x int he first equation x + 2y = 9 (y + 6) + 2y = 9 3y + 6 = 9 3y = 3 y = 1

hero (hero):

x + 2y = 9 => x = 9 - 2y x - y = 6 => x = 6 + y x = x 9 - 2y = 6 + y 9-6 = y + 2y 3 = 3y 3/3 = y 1 = y 7 = x

hero (hero):

So the correct point is (x,y) = (7,1)